Please be as clear as possible.

Textbook - Applied Statistics and Probability for Engineers by Montgomery, 6th EditionAn engineer performed N= 20 tests to assess the load capacity of a new anchoring device. The measured load capacities are as
Answer #1

Final answers are highlighted in colour.

Q1.

a)Calculated using Excel.

Mean(\mu)= AVERAGE(A1:A20)= 19.1

Standard Deviation(\sigma)= STDEV.P(A1:A20)= 6.402

Coefficient of variation= \frac{\sigma}{\mu} = 6.402/19.1= 0.3352= 33.52%

Q2.

b) Mean(\overline{X})= AVERAGE(C1:C9)= 17

Standard Deviation(S)= 5.809

Coefficient of variation= \frac{\sigma}{\mu} = 5.809/17= 0.3417= 34.17%

c) Standard Error on the mean(SE)= S Vn =5.809/sqrt(9)= 5.809/3= 1.9365

d)90% Confidence interval

= \overline{X}\pm t^*SE

t* for 90% confidence interval= 1.859547

CI= \17\pm 1.859547 * 1.9365 = (13.399, 20.601)

Q3.

e)

Expert's Hypothesis

H0: \mu = 20

Ha: \mu \neq 20

Sample mean(\overline{X})=17, Standard Error(SE)=1.9365

t stat= \frac{\overline{X}-\mu}{SE} = \frac{17-20}{1.9365} = -1.549

p value for t= -1.549, degrees of freedom= 8 is 0.159974

As p value(0.159974) > level of significance(0.05), we cant reject the null hypothesis.

There is no sufficient evidence to say that the mean value is different than 20.

Q4.

f)

For the new sample,

Mean(\overline{X})= 21.11

Standard Deviation(S)= 7.59

Coefficient of variation= \frac{\sigma}{\mu} = 7.59/21.11= 0.3595= 35.95%

Q5.

a)Length of the confidence interval should not exceed 2.

t^* \frac{S}{\sqrt{n}} \leq 2

If t* is assumed to be a constant.

\frac{1.859547*5.809}{\sqrt{n}} \leq 2

Solving for n gives, n>29

20 more tests should be done

But,

As degrees of freedom also changes with sample size(n), t* also varies. So this needs to be performed on Solver to get the value of n.

Solving the equation in solver gives n>= 25